Calcium has a work function of 2.7 eV. (a) What is the function of calcium expressed in joules? (b) What is the threshold frequency for calcium? (c) What is the maximum wavelength that will cause emission from calcium?
(e = 1.6 x 10-19 c,h - 6.6 x 10-34 Js,c = 3.0 x 108 ms-l.)
1
Expert's answer
2020-11-10T07:07:57-0500

Solution

a) changing work function into joules

"\phi=2.7\times 1.6\times10^{-19}"

"\phi=4.32\times10^{-19}J"

b) for threshold frequency

"\phi=h\nu_0"

"\nu_0=\frac{\phi}{h}=\frac{4.32\times10^{-19}}{6.6\times10^{-34}}=65.45\times10^{-15} Hz"

C) now maximum wavelength

"\lambda=\frac{c}{\nu_0}=\frac{3\times10^8}{65.45\times10^{15}}=45.83 \times 10^{-10}m"